An isosceles triangle has an angle that measures 80°. Which other angles could be in that isosceles triangle? Choose all that apply.
Found 3 solutions by dkppathak, MathTherapy, Edwin McCravy:
Answer by dkppathak(439)

An isosceles triangle has an angle that measures 80°. Which other angles could be in that isosceles triangle? Choose all that apply.
solution
in isosceles triangle two angle will be equal and sum of all angle will be 180 degree
let equal angle are x each
tringle property
x+x+80=180
2x=180-80=100
x=100/2=50 degree
angles are 50,50,80 degree

Answer by MathTherapy(10556)

An isosceles triangle has an angle that measures 80°. Which other angles could be in that isosceles triangle? Choose all that apply.
Correct answers:  


Answer by Edwin McCravy(20060)

He forgot one of the possible angles.  He just gave you 50°, and put 80° for the
other one, but that was given. Here are the two isosceles triangles you could
have.  What is the skinny angle at the top of the skinny triangle on the left?



Edwin